# Efficient Reinforcement Finetuning via Adaptive Curriculum Learning

### v2

@@ 5.1 Training Efficiency
As shown in Figure 3 and Table 1, models trained with AdaRFT consistently require fewer training steps to match the performance of those trained with standard PPO or PPO on filtered data. Specifically, we report how many additional steps are needed for PPO variants to match the performance of AdaRFT at step 60 for Qwen 2.5 Math 1.5B, and step 40 for Qwen 2.5 7B. Because models are evaluated only every 5 training steps, we apply exponential smoothing with a smoothing parameter to the accuracy curves to reduce variance and obtain stable estimates of performance over time. The shaded areas in Figure 3 represent the raw, unsmoothed accuracy 1%, offering a visual cue for the typical fluctuation in evaluation accuracy.
| Model | Setup | Method | Avg Step Time (s) | Extra Steps (%) |
| Qwen2.5 Math 1.5B | skew-difficult | AdaRFT | 122.24 | +0 (0.0%) |
| PPO | 132.95 | +43 (71.7%) | ||
| PPO (w/ Filter) | 128.20 | +49 (81.7%) | ||
| uniform | AdaRFT | 121.31 | +0 (0.0%) | |
| PPO | 126.82 | +34 (56.7%) | ||
| PPO (w/ Filter) | 126.35 | +52 (86.7%) | ||
| skew-easy | AdaRFT | 120.52 | +0 (0.0%) | |
| PPO | 121.15 | +16 (26.7%) | ||
| PPO (w/ Filter) | 115.12 | +21 (35.0%) | ||
| Qwen2.5 7B | skew-difficult | AdaRFT | 239.92 | +0 (0.0%) |
| PPO | 246.21 | +24 (60.0%) | ||
| PPO (w/ Filter) | 254.22 | +25 (62.5%) | ||
| uniform | AdaRFT | 234.16 | +0 (0.0%) | |
| PPO | 243.82 | +13 (32.5%) | ||
| PPO (w/ Filter) | 263.11 | +23 (57.5%) | ||
| skew-easy | AdaRFT | 247.44 | +0 (0.0%) | |
| PPO | 235.27 | +20 (50.0%) | ||
| PPO (w/ Filter) | 233.13 | +17 (42.5%) |
For Qwen 2.5 Math 1.5B, standard PPO requires 43 extra steps (+71.7%) in the skew-difficult setting and 34 steps (+56.7%) in the uniform setting to match AdaRFT’s performance. PPO with filtered training data requires even more: +49 steps (81.7%) and +52 steps (86.7%) in the respective settings. In the skew-easy scenario, PPO requires +16 steps (26.7%), while PPO with filtered data needs +21 steps (35.0%) to catch up to AdaRFT. The efficiency gains remain significant with the larger Qwen 2.5 7B model. In the skew-difficult setting, PPO and PPO with filtered data require +24 steps (60.0%) and +25 steps (62.5%), respectively.
In addition to improved sample efficiency, AdaRFT also achieves faster average training time per step across nearly all settings, as reported in Table 1. Though PPO with filtered data can sometimes offer marginal gains in per-step time (e.g., skew-easy setups), it still falls behind in convergence speed. While the underlying causes are discussed in Section 6, these results underscore that AdaRFT is both sample-efficient and compute-efficient, delivering faster and more cost-effective training.
@@ 5.2 Model Performance
In addition to improving efficiency, AdaRFT (PPO) also improves the final model performance. As shown in Table 2, at the end of training (step 100), AdaRFT yields consistent improvements in final accuracy across all configurations. The reported averages reflect accuracy across six diverse benchmarks: GSM8K, MATH 500, OlympiadBench, Minerva Math, AMC 23, and AIME 24.
@@
On skew-difficult data, the Qwen 2.5 Math 1.5B model improves from 37.41% (PPO) to 40.48% (AdaRFT (PPO)), a gain of over 3 percentage points in average accuracy. Similar improvements appear in the uniform setting, where AdaRFT (PPO) reaches 41.11%, compared to 37.20% with PPO. Even on skew-easy data, where the baseline performs well, curriculum learning still improves performance, reaching 39.18% versus 38.46%. For the larger Qwen 2.5 7B model, final accuracy gains are also consistent, though slightly more modest. In the skew-difficult setting, AdaRFT (PPO) improves from 44.17% to 46.83%. In the uniform setting, accuracy rises from 44.70% to 46.92%, and in the skew-easy case, from 45.07% to 45.94%. These results show that AdaRFT is effective even for stronger models, enhancing both stability and peak performance.
| Model | Setup | Method |
Average | |||||||||||
| Qwen 2.5 Math 1.5B | skew-difficult | PPO | 69.67 | 64.60 | 20.65 | 12.87 | 47.50 | 9.17 | 37.41 | |||||||||||
| AdaRFT | 74.00 | 66.40 | 20.36 | 15.07 | 55.00 | 12.08 | 40.48 | |||||||||||||
| uniform | PPO | 71.95 | 65.20 | 21.10 | 15.81 | 42.50 | 6.67 | 37.20 | ||||||||||||
| AdaRFT | 74.53 | 66.20 | 21.99 | 14.34 | 57.50 | 12.08 | 41.11 | |||||||||||||
| skew-easy | PPO | 72.71 | 67.40 | 19.17 | 13.97 | 45.00 | 12.50 | 38.46 | ||||||||||||
| AdaRFT | 73.62 | 66.20 | 19.91 | 13.97 | 55.00 | 9.17 | 39.18 | |||||||||||||
| Qwen 2.5 7B | skew-difficult | PPO | 89.69 | 71.20 | 23.33 | 23.53 | 50.00 | 11.25 | 44.17 | |||||||||||
| AdaRFT | 90.98 | 71.40 | 25.85 | 22.43 | 52.50 | 15.83 | 46.83 | |||||||||||||
| uniform | PPO | 89.31 | 72.40 | 23.63 | 25.37 | 42.50 | 15.00 | 44.70 | ||||||||||||
| AdaRFT | 90.14 | 72.60 | 24.96 | 24.26 | 55.00 | 14.58 | 46.92 | |||||||||||||
| skew-easy | PPO | 89.39 | 73.60 | 23.33 | 24.26 | 47.50 | 13.33 | 45.07 | ||||||||||||
| AdaRFT | 90.14 | 72.60 | 25.56 | 23.16 | 50.00 | 14.17 | 45.94 |
6 Discussion
6.1 Analysis on Efficiency and Accuracy
Sample efficiency and final performance.
AdaRFT substantially improves both sample efficiency and final model performance, particularly in non-uniform data distributions. By dynamically adjusting task difficulty during training, AdaRFT enables models to allocate learning capacity more effectively, focusing on simpler tasks early on and gradually transitioning to harder ones. This leads to faster convergence and better use of training steps. The performance gains are especially pronounced in the Qwen 2.5 Math 1.5B model, where curriculum structure plays a larger role in guiding learning. For the larger Qwen 2.5 7B model, improvements are still evident, though more modest—likely because the model’s increased capacity allows it to handle diverse and imbalanced tasks without as much external scaffolding.
Reduced training time per step.
AdaRFT also results in a slightly lower average training time per step. This is largely due to the fact that easier problems require fewer tokens to solve. For example, an arithmetic reasoning question from GSM8K might require only around 200 tokens for Qwen 2.5 7B to reach a correct answer, whereas a competition-level math problem from AIME could require around 2000 tokens—a 10 difference in rollout length. The total token length affects multiple components of the training step, including the rollout itself and the subsequent PPO update. While PPO update time does not scale linearly with sequence length due to batching and attention computation patterns, longer sequences still incur higher compute costs. As a result, curriculum learning’s tendency to prioritize shorter, easier problems early in training leads to shorter sequences on average, reducing per-step compute and improving overall training throughput.
When does curriculum learning help?
Our findings show that curriculum learning provides the greatest benefits under two key conditions: (1) imbalanced training distributions, and (2) limited model capacity. In skewed distributions, particularly the skew-difficulty settings, standard PPO often struggles to gain traction early in training due to insufficient reward signals. AdaRFT mitigates this by initially sampling easier problems, enabling the model to bootstrap capabilities before tackling harder content. For example, in the Qwen 2.5 Math 1.5B model trained on skew-difficult data, AdaRFT improves final accuracy by over 3 percentage points and reduces training steps by 44%. This suggests that AdaRFT is especially helpful when the model would otherwise be under-optimized due to overly difficult or unbalanced data.
Conversely, the benefits of AdaRFT are less pronounced when the model is strong enough or the data is already well-balanced. For instance, in the uniform setup with the Qwen 2.5 7B model, AdaRFT offers only modest improvements in convergence speed and final accuracy. Likewise, in the skew-easy setup for the smaller Qwen 2.5 Math 1.5B model, the curriculum yields relatively smaller gains. In both cases, the model is either already exposed to a representative distribution of task difficulties or finds most problems challenging enough, thus reducing the need for dynamic difficulty adjustment.
@@ 6.1 Analysis on Efficiency and Accuracy
It is important to note that manual data curation and task scheduling, carefully designed for a specific model, could potentially achieve similar results by selecting a training sequence that aligns with the model’s learning capacity (Yu et al., 2025; Shen et al., 2025; Chen et al., 2025; Zeng et al., 2025). However, such methods require significant human effort, domain knowledge, and often need to be re-tuned for each new model or task. In contrast, AdaRFT requires no manual data curation or model-specific preprocessing. The curriculum automatically adapts to the model’s reward signal during training, making it broadly applicable across model scales and training distributions. This automatic adaptability not only saves engineering effort but also improves scalability and robustness in real-world training pipelines. Moreover, AdaRFT is particularly advantageous in fixed data settings, as it can tailor the difficulty schedule to match the capabilities of any model, whether strong or weak, without altering the dataset, providing a unified solution that generalizes across model architectures and skill levels.
6.2 Data Difficulty on Model Performance
@@
6.3 Ablation on Target Reward
6.4 Training on LLM-Estimated Difficulty
@@
6.5 AdaRFT with Diverse RL Algorithms
To evaluate the generality of AdaRFT beyond PPO, we trained the Qwen 2.5 Math 1.5B model on a skew-difficult data distribution using two alternative reinforcement learning algorithms: REINFORCE++ and GRPO (see implementation details in Appendix A). As shown in Figure 7, AdaRFT significantly improves both the convergence speed and final accuracy across these variants. Across both cases, the adaptive curriculum acts orthogonally to the underlying optimization method. These results reinforce the plug-and-play nature of AdaRFT: it consistently enhances sample efficiency and policy robustness across algorithmic choices, making it broadly applicable in diverse reinforcement finetuning pipelines. Notably, this generalization holds without any additional tuning or algorithm-specific modifications, underscoring the practical utility of curriculum-aware training in both lightweight and computation-heavy RFT settings.
7 Conclusion
Acknowledgement
The authors would like to thank the members of the LIME lab and the USC NLP group for their helpful insights and feedback.
Appendix A Implementation Details
A.1 Training Configuration

### v3

@@ 5.1 Training Efficiency
As shown in Figure 3 and Table 1, models trained with AdaRFT consistently require fewer training steps to match the performance of those trained with standard PPO, PPO on filtered data, and PPO with a fixed curriculum schedule. Specifically, we report how many additional steps are needed for PPO variants to match the performance of AdaRFT at step 60 for Qwen 2.5 Math 1.5B, and step 40 for Qwen 2.5 7B. Because models are evaluated only every 5 training steps, we apply exponential smoothing with a smoothing parameter of to the accuracy curves to reduce variance. The shaded areas in Figure 3 represent the raw, unsmoothed accuracy 1%, offering a visual cue for the typical fluctuation in evaluation accuracy. For Qwen 2.5 Math 1.5B, standard PPO requires 43 extra steps (+71.7%) in the skew-difficult setting and 34 steps (+56.7%) in the uniform setting to match AdaRFT’s performance. PPO with filtered training data requires even more: +49 steps (81.7%) and +52 steps (86.7%) in the respective settings. In the skew-easy scenario, PPO requires +16 steps (26.7%), while PPO with filtered data needs +21 steps (35.0%) to catch up to AdaRFT. The efficiency gains remain significant with the larger Qwen 2.5 7B model. In the skew-difficult setting, PPO and PPO with filtered data require +24 steps (60.0%) and +25 steps (62.5%), respectively. PPO with a fixed curriculum schedule also follows this trend, suggesting that while fixed curricula can modestly improve training efficiency, their inability to adapt the the model’s evolving learning dynamics limits their convergence speed relative to AdaRFT.
In addition to improved sample efficiency, AdaRFT also achieves faster average training time per step across nearly all settings, as reported in Table 1. This is largely due to the fact that easier problems require fewer tokens to solve. For example, an arithmetic reasoning question from GSM8K might require only around 200 tokens to reach a correct answer, whereas a competition-level math problem from AIME could require around 2000 tokens, a 10 difference in rollout length. The total token length affects multiple components of the training step, including the rollout itself and the subsequent PPO update. While PPO update time does not scale linearly with sequence length due to batching and attention computation patterns, longer sequences still incur higher compute costs. As a result, curriculum learning’s tendency to prioritize easier problems early in training leads to shorter sequences on average, reducing per-step compute and improving overall training throughput.
@@ 5.2 Model Performance
While the primary goal of AdaRFT is to improve training efficiency, it also consistently maintains and often improves final model performance. As shown in Tables 1, 2, and Figure 3, AdaRFT achieves higher final accuracy at the end of training (step 100) across all configurations. Reported results are averaged over six benchmarks: GSM8K, MATH 500, OlympiadBench, Minerva Math, AMC 23, and AIME 24. On skew-difficult data, Qwen 2.5 Math 1.5B improves from 37.41% with PPO to 40.48% with AdaRFT, a gain of over 3 percentage points. Similar gains are observed in the uniform setting (41.11% vs. 37.20%). Even on skew-easy data, where the baseline already performs well, AdaRFT still yields modest improvements (39.18% vs. 38.46%). For the larger Qwen 2.5 7B model, improvements are smaller but consistent: from 44.17% to 46.83% on skew-difficult data, from 44.70% to 46.92% in the uniform setting, and from 45.07% to 45.94% on skew-easy data. These results demonstrate that AdaRFT improves training efficiency without sacrificing final accuracy, and in practice often enhances peak performance. A more detailed per-benchmark breakdown is provided in Appendix A.6.
6 Conclusion
Limitations
When Does Curriculum Learning Help Most?
Our findings show that curriculum learning provides the greatest benefits under two key conditions: (1) imbalanced training distributions, and (2) limited model capacity. In skewed distributions, particularly the skew-difficult settings, standard PPO often struggles to gain traction early in training due to insufficient reward signals. AdaRFT mitigates this by initially sampling easier problems, enabling the model to bootstrap capabilities before tackling harder content. Conversely, the benefits of AdaRFT are less pronounced when the model is strong enough or the data is already well-balanced. In both cases, the model is either already exposed to a representative distribution of task difficulties or finds most problems challenging enough, thus reducing the need for dynamic difficulty adjustment. In addition, we conducted further experiments detailed in Appendix A, including ablation studies on different target reward (A.2), evaluations on datasets with more extreme difficulty distributions (A.3), difficulty estimation using an LLM-based judge (A.5), and instantiations of AdaRFT with alternative RL algorithms (GRPO, REINFORCE++) (A.4). Across all these settings, AdaRFT consistently demonstrates effectiveness, highlighting its robustness to diverse data distributions, compatibility with various RL algorithms, and flexibility with different difficulty metrics.
@@ 6 Conclusion
It is important to note that manual data curation and task scheduling carefully designed for a specific model could potentially achieve similar results by selecting a training sequence that aligns with the model’s learning capacity (Yu et al., 2025; Shen et al., 2025; Chen et al., 2025; Zeng et al., 2025). However, such methods require significant human effort, domain knowledge, and often need to be re-tuned for each new model or task. Fixed curriculum schedules face similar limitations: because they rely on a predetermined ordering of difficulty, they cannot adjust when the model learns faster or slower than expected. In contrast, AdaRFT requires no manual data curation or model-specific preprocessing. The curriculum automatically adapts to the model’s reward signal during training, making it broadly applicable across model scales and training distributions. This automatic adaptability not only saves engineering effort but also improves scalability and robustness in real-world training pipelines. Moreover, AdaRFT is particularly advantageous in fixed data settings, as it can tailor the difficulty schedule to match the capabilities of any model, whether strong or weak, without altering the dataset, providing a unified solution that generalizes across model architectures and skill levels.
Use of Offline Difficulty Estimation.
A core design choice of AdaRFT is to rely on offline difficulty estimates that remain fixed throughout training, rather than recomputing difficulty online as the model improves. AdaRFT adopts a solver-independent notion of difficulty, in which problems such as IMO-level questions are considered inherently more difficult than grade-school math problems and are assigned a single difficulty score based on intrinsic complexity, in contrast to online curriculum methods that use solver-dependent definitions that evolve with the model’s success rate. This design aligns naturally with many mathematics competitions (Hendrycks et al., 2021) and puzzle datasets (Lin et al., 2025), which provide fixed difficulty annotations, and with modern RL fine-tuning pipelines where pass@ rollouts are commonly computed during data preparation regardless of curriculum learning. Recent systems such as Open Reasoner Zero (Hu et al., 2025), Kimi k1.5 (Team et al., 2025), and ZR1 1.5B (Zyphra, 2025) evaluate pass@ on the full dataset prior to training to filter overly easy or difficult examples, making difficulty information readily available in practice. Rather than updating difficulty estimates themselves, AdaRFT adapts the curriculum by dynamically updating the target difficulty based on the model’s success rate on the currently sampled data, achieving the same functional outcome as online difficulty estimation while remaining more stable, efficient, and scalable for large-scale RL fine-tuning.
Appendix A Further Discussion
A.1 Training Dynamics: AdaRFT vs. Fixed Curriculum vs. PPO
To further contextualize the role of adaptive curricula, we analyze the training dynamics of fixed curriculum (PPO w/ FC), standard PPO, and AdaRFT when training the Qwen 2.5 Math 1.5B model on the uniform distribution. Figure 4 illustrates the evolution of accuracy, reward, response length, and sampled difficulty across the first 100 training steps.
The accuracy curves highlight the central tradeoff of fixed curricula. PPO w/ FC initially converges slightly faster than standard PPO, benefiting from early exposure to easier problems. However, it ultimately underperforms compared to AdaRFT because its difficulty schedule increases independently of the model’s actual learning progress. As a result, the model is pushed into harder problem regimes too quickly. This misalignment is clearly visible in the reward curve. Early in training, PPO w/ FC achieves rewards well above 0.5, indicating that the model is initially exposed to problems that are easier than its current capability. However, as training progresses, the fixed curriculum increases difficulty at a rate that outpaces the model’s learning speed. As a result, the reward drops below 0.5 and continues declining, showing that the model is increasingly confronted with problems it cannot yet solve. This mismatch limits the effectiveness of the updates and ultimately leads to slower convergence compared to an adaptive curriculum.
The response length patterns reinforce this interpretation. Because longer responses typically correspond to more difficult reasoning tasks, the rapid increase in response length under PPO w/ FC shows that the curriculum escalates difficulty faster than the model can adapt. In contrast, standard PPO maintains more stable lengths but lacks the structured progression necessary for efficient learning. AdaRFT, by comparison, keeps response lengths moderate and gradually increasing, consistent with its difficulty traces: the curriculum raises problem difficulty only when the model’s reward stays near the target value, ensuring that the model always receives examples that are challenging but solvable.
This particular dynamic suggest a general underlying issue: since the model’s perceived difficulty changes over the course of training, a fixed curriculum cannot remain aligned, and different model-dataset combinations may experience extended periods of being over-challenged or under-challenged. This mismatch limits the effectiveness of updates and leads to slower convergence compared to an adaptive curriculum. In contrast, AdaRFT is designed to handle precisely this challenge: by adjusting the difficulty schedule based on the model’s reward signal, it continually matches training difficulty to the model’s evolving capability, ensuring sustained learning progress throughout training. Without the ability to adjust to real-time model performance, fixed schedules risk either overwhelming the model or wasting compute on overly easy tasks. AdaRFT avoids both extremes by maintaining the average reward near 0.5, automatically pacing the introduction of harder problems as the model improves. This adaptive alignment between task difficulty and model capability leads to smoother reward trajectories, controlled response lengths, and ultimately more stable and efficient learning.
A.2 Ablation on Target Reward
A.3 Data Difficulty on Model Performance
@@
A.4 AdaRFT with Diverse RL Algorithms
To evaluate the generality of AdaRFT beyond PPO, we trained the Qwen 2.5 Math 1.5B model on a skew-difficult data distribution using two alternative reinforcement learning algorithms: REINFORCE++ and GRPO (see implementation details in Appendix B). As shown in Figure 7, AdaRFT significantly improves both the convergence speed and final accuracy across these variants. Across both cases, the adaptive curriculum acts orthogonally to the underlying optimization method. These results reinforce the plug-and-play nature of AdaRFT: it consistently enhances sample efficiency and policy robustness across algorithmic choices, making it broadly applicable in diverse reinforcement finetuning pipelines. Notably, this generalization holds without any additional tuning or algorithm-specific modifications, underscoring the practical utility of curriculum-aware training in both lightweight and computation-heavy RFT settings.
A.5 Training on LLM-Estimated Difficulty
@@
A.6 Detailed Breakdown of Model Performance
Table 2 provides a per-benchmark breakdown of final accuracy at step 100 across all models, data setups, and training methods. Results are reported on six diverse mathematical reasoning benchmarks spanning grade-school arithmetic (GSM8K), competition-style problems (MATH 500, AMC 23, AIME 24), and advanced proof-oriented datasets (OlympiadBench, Minerva Math). Across both model scales and all data distributions, AdaRFT either matches or outperforms PPO-based baselines on the majority of individual benchmarks, resulting in consistent gains in the overall average accuracy.
For Qwen 2.5 Math 1.5B, improvements from AdaRFT are particularly pronounced on GSM8K, MATH 500, AMC 23, and AIME 24 under skew-difficult and uniform settings, indicating stronger generalization across both structured and competition-style problems. Even in the skew-easy setting, where baseline PPO already performs well, AdaRFT maintains or slightly improves performance without introducing regressions on harder benchmarks such as OlympiadBench and Minerva Math. For the larger Qwen 2.5 7B model, gains are more modest but remain consistent, with AdaRFT achieving higher average accuracy across all three data distributions and stable performance on the most challenging benchmarks. Overall, these results confirm that the efficiency gains of AdaRFT do not come at the cost of final model quality, and in practice often translate into improved performance across a broad range of mathematical reasoning tasks.
| Model | Setup | Method |
Average | |||||||||||
| Qwen 2.5 Math 1.5B | skew-difficult | AdaRFT | 74.00 | 66.40 | 20.36 | 15.07 | 55.00 | 12.08 | 40.48 | |||||||||||
| PPO | 69.67 | 64.60 | 20.65 | 12.87 | 47.50 | 9.17 | 37.41 | |||||||||||||
| PPO (w/ FC) | 72.55 | 66.40 | 20.95 | 14.34 | 45.00 | 4.06 | 37.22 | |||||||||||||
| uniform | AdaRFT | 74.53 | 66.20 | 21.99 | 14.34 | 57.50 | 12.08 | 41.11 | ||||||||||||
| PPO | 71.95 | 65.20 | 21.10 | 15.81 | 42.50 | 6.67 | 37.20 | |||||||||||||
| PPO (w/ FC) | 71.95 | 65.60 | 19.61 | 14.71 | 42.50 | 9.27 | 37.27 | |||||||||||||
| skew-easy | AdaRFT | 73.62 | 66.20 | 19.91 | 13.97 | 55.00 | 9.17 | 39.18 | ||||||||||||
| PPO | 72.71 | 67.40 | 19.17 | 13.97 | 45.00 | 12.50 | 38.46 | |||||||||||||
| PPO (w/ FC) | 72.40 | 66.20 | 20.06 | 13.24 | 50.00 | 6.67 | 38.09 | |||||||||||||
| Qwen 2.5 7B | skew-difficult | AdaRFT | 90.98 | 71.40 | 25.85 | 22.43 | 52.50 | 15.83 | 46.83 | |||||||||||
| PPO | 89.69 | 71.20 | 23.33 | 23.53 | 50.00 | 11.25 | 44.17 | |||||||||||||
| PPO (w/ FC) | 89.92 | 72.00 | 24.52 | 25.37 | 47.50 | 13.96 | 45.54 | |||||||||||||
| uniform | AdaRFT | 90.14 | 72.60 | 24.96 | 24.26 | 55.00 | 14.58 | 46.92 | ||||||||||||
| PPO | 89.31 | 72.40 | 23.63 | 25.37 | 42.50 | 15.00 | 44.70 | |||||||||||||
| PPO (w/ FC) | 89.84 | 72.40 | 23.92 | 23.90 | 42.50 | 13.33 | 44.32 | |||||||||||||
| skew-easy | AdaRFT | 90.14 | 72.60 | 25.56 | 23.16 | 50.00 | 14.17 | 45.94 | ||||||||||||
| PPO | 89.39 | 73.60 | 23.33 | 24.26 | 47.50 | 13.33 | 45.07 | |||||||||||||
| PPO (w/ FC) | 89.46 | 72.20 | 24.37 | 23.90 | 37.50 | 13.33 | 43.46 |
Appendix B Implementation Details
B.1 Training Configuration